J&W Corporation manufactures a new electronic game console. The current standard cost sheet for a game console follows:

Direct materials, ? kilograms at $6 per kilogram $ ? per game
Direct labor, 0.75 hours at ? per hour ? per game
Overhead, 0.75 hours at ? per hour ? per game
Total costs $ 39 per game

Assume that the following data appeared in J&W’s records at the end of the past month:

Actual production 58,000 units
Actual sales 55,000 units
Materials (132,000 kilograms) $ 851,000
Materials price variance 59,000 U
Materials efficiency variance 43,800 U
Direct labor price variance 32,800 U
Direct labor (41,000 hours) 754,400
Underapplied overhead (total) 20,000 U

There are no materials inventories.

Required:

a. Complete the standard cost sheet for a game console given below. Compute the direct labor efficiency variance. (Do not round intermediate calculations. Round your answers to 2 decimal places. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)

Direct materials kilograms at $6 per kilogram per game
Direct labor, 0.75 hours at per hour per game
Overhead, 0.75 hours at per hour per game
Total costs $39 per game
Direct labor:
Efficiency variance   

b. Assume that all production overhead is fixed and that the $20,000 underapplied is the only overhead variance that can be computed. What are the actual and applied overhead amounts?

Overhead
Actual
Applied
Answer #1

a.

Direct materials $2.15 kilograms at $6 per kilogram 12.9 per game
Direct labor, 0.75 hours at $17.60 per hour 13.2 per game
Overhead, 0.75 hours at $17.20 per hour 12.9 per game
Total costs $39 per game
Direct labor:
Efficiency variance $44,000 F

Working:

Material efficiency variance = $43,800 U
Actual price of material = $851,000 / 132000 kd = $6.44696 per kg
Actual quantity of material = 132000 Kg
Standard price of material = $6 per kg
(SQ - AQ) * SP = -$43,800
(SQ - 132000) * $6 = -$43,800
SQ = 124700 Kg
Actual production = 58000 units
Standard quantity of material per Kg = 124700 / 58000 = 2.15 Kg
Direct labor incurred = $754,400
Actual hours of direct labor = 41000
Actual rate of direct labor = $754,400 / 41000 = $18.40 per hour
Direct labor price variance = $32,800 U
(SR - AR)* AH = -$32,800
(SR - $18.40) * 41000 = -$32,800
SR = $17.60 per hour
Direct labor efficiency variance = (SH - AH) * SR = (43500 - 41000) * $17.60 = $44,000 F
Standard hours for actual production = 58000 * 0.75 = 43500 hours

b.

Overhead
Actual $    768,200.00
Applied $    748,200.00

Working:

Overhead applied = 58000*0.75*$17.20 = $748,200
Underapplied overhead = $20,000 U
Actual applied overhead = Overhead applied + Underapplied overhead = $748,200 + $20,000 = $768,200